SpaceRef

SpaceRef


SATOPS Morning Report - March 14, 2003 Day 073

Status Report From: China National Space Administration
Posted: Friday, March 14, 2003

DMSP

DMSP operations were nominal over the past 24 hours.

POES

NOAA-15 rev 25111 / F at 1908Z on March 13: Executed SCR #62, to run CP TX1SS. This command procedure was used to make changes to the STESM parameters for N-15. Commanding was uneventful and all systems appear nominal.

POES operations were nominal over the past 24 hours.

GOES

At 0524Z on March 14, a GOES-10 frame break was observed and attributed to high BER at CDA (unknown cause).

Eclipse operations for all spacecraft continue. Eclipse schedules for GOES-8 and GOES-10 (which can be used to track the day-to-day image deletions due to KOZ and eclipse) can be found at: http://www.oso.noaa.gov/goes/koz-eclipse/koz-eclipse.htm

GOES-12 remains in a test phase, in normal on-orbit mode, drifting at an eastward rate of ~0.33 degrees per day toward the GOES-East position where it will replace GOES-8 in April. Routine SXI imaging is proceeding along with SXI test exposures. Between 2002Z and 2132Z, magnetometer testing was performed as per SCP.

GOES-9 is currently at approximately 174.3 W drifting at a westward rate of ~0.735 degrees per day toward 155 E longitude where it will support the GMS-5 replacement mission in April. Canberra and Goldstone DSN ranging activities are continuing.

GOES operations were nominal over the past 24 hours.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter